OpenStudy (anonymous):

A meterstick standing vertically casts a shadow 1.5 meters long on the ground. At the same time, a nearby tree casts a shadow 9 meters long on the ground. About how tall is the tree?

OpenStudy (anonymous):



OpenStudy (anonymous):

Use similar triangles. Call x = tree's height. x/1 = 9/1.5 --> x = ....